OpenStudy (anonymous):

Mathematical Induction Question: Prove using MI : ( x^n - y^n) is divisible by (x-y)

OpenStudy (nikvist):

\[x^{n+1}-y^{n+1}=x^{n+1}-xy^n+xy^n-y^{n+1}+x^ny-x^ny=\] \[=x(x^n-y^n)+y(x^n-y^n)-xy(x^{n-1}-y^{n-1})=\] \[=(x+y)(x^n-y^n)-xy(x^{n-1}-y^{n-1})\] this is divisible by x-y

OpenStudy (anonymous):

How is this divisible by x-y ?